OpenStudy (anonymous):

log[4] (x+2) - log[4] (x-1) =1

Nnesha (nnesha):

\[\huge\rm log_4 (x+2) \color{red}{-} \log_4 (x-1) =1\] apply quotient property \[\huge\rm log_b x \color{red}{-} \log_b y = \log_b \frac{ x }{ y }\] if there is negative sign you can chang3 that to divison subtraction ---->division addition .-----> multiplication

OpenStudy (anonymous):

yes i understand addition becomes multiplication and so on

Nnesha (nnesha):

so who would you write it \[\huge\rm log_4 (x+2) \color{red}{-} \log_4 (x-1) =1\] ?

OpenStudy (anonymous):

That's the part that is confusing me

Nnesha (nnesha):

both log have same base which is 4 so change \[\huge\rm log_4 (x+2) \color{red}{-} \log_4 (x-1) \] this to division like in this example \[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y }\]

OpenStudy (anonymous):

\[\log[4]\frac{ x+2 }{ x-1 } = 1 ?\]

Nnesha (nnesha):

yep right now change log to exponential form \[\huge\rm log = \log_{10}\] change log to exponential form for example \[\huge\rm log_b x = y \]\[\huge\rm b^y = x\]